OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5). 10.

OpenStudy (anonymous):

paralel, means with same slope. Slope of x+4y=6 is -1/4, because: y=-1/4x+3/2 so your line is: (y-5)=-1/4(x+8) y=-1/4x+3